OpenStudy (anonymous):

3) If sin(theta+36)=cos theta, find theta

OpenStudy (a_clan):

\[\sin(\Theta +36) = \cos (\Theta)\] \[\sin(\Theta +36) =\sin (90-\Theta)\] So, \[\theta +36 = 90 -\theta\]

OpenStudy (anonymous):

θ=27